OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

x = 3 / 8

OpenStudy (anonymous):

that isnt an option though

OpenStudy (anonymous):

2x + y = 1 4x + 2y = −1 There is no x value as there is no solution.

OpenStudy (anonymous):

what are the options then ?

OpenStudy (anonymous):

it is either 1.x = 0 2. x = 2 3. There is no x value as there is no solution. 4. x can be any value as there is an infinite number of solutions.

OpenStudy (jdoe0001):

can you solve for "y" on the 1st equation? what would that give you?

OpenStudy (anonymous):

it is 3 like I said.